OpenStudy (anonymous):

Given the two ordered pairs, find the slope of the line that passes through them. (-8, 9) (3, -5)

OpenStudy (saifoo.khan):

Do you know the slope formula?

OpenStudy (anonymous):

@gitahimart

OpenStudy (anonymous):

use the slope formula: slope = (y1-y2)/(x1-x2) =(9--5)/(-8-3) =??

OpenStudy (anonymous):

\[m~= ~~{y_{2}-y_{1}\over x_{2}-x_{1}}~~~\rightarrow~~{-5-9\over3-(-8)}~~~\rightarrow~~{-14\over11}\]